OpenStudy (anonymous):

In pentagon ABCDF, the measures of angle A and angle B are both One hundred degrees The measures of angle C and angle D are both One hundred fifteen degrees What is the measure of angle F? seventy degrees one hundred ten degrees four hundred seventy degrees three hundred twenty five degrees

OpenStudy (aivantettet26):

could you draw it?

OpenStudy (anonymous):

Are your options: A. 70º seventy degrees B. 110º one hundred ten degrees C. 470º four hundred seventy degrees D. 325º three hundred twenty five degrees

OpenStudy (anonymous):

For a pentagon: S(i) = 180x(5-2) = 180x3 = 540 degrees

OpenStudy (anonymous):

yes

OpenStudy (radar):

Oh I see you know they sum up to 540 degrees. That is the key.

OpenStudy (radar):

Sum up what you have given by the problem.

OpenStudy (anonymous):

Really you just need to subtract now. \[540−(100+100+115+115)=? \]\[540-200-230= 110\] So your answer is B!